# GLOBALIZATION? TRADE WAR? A COUNTERBALANCE PERSPECTIVE

### v4

@@ 7 Discussions
We let $P_{ji}$ have a small change $\Delta P_{ji}$ and attempt to calculate its effect on $\pi$, including $\pi_{i}$, $\pi_{j}$, and $\pi_{-i}$. Accordingly, $P$ have three other simultaneous changes: $-\Delta P_{ji}$ on $P_{jj}$ to offset the change on $P_{ji}$ in the same row; $\lambda_{ji}\Delta P_{ji}$ on $P_{ij}$ for country $j$’s retaliation upon $i$’s change at $P_{ji}$; and $-\lambda_{ji}\Delta P_{ji}$ on $P_{ii}$ to maintain the unit sum of the $i$th row. Therefore, $\frac{\mathrm{d}P}{\mathrm{d}P_{ji}}=\lim\limits_{\Delta P_{ji}\to 0}\frac{\Delta P}{\Delta P_{ji}}$ is a zero $n\times n$ matrix except for: $1$ at $(j,i)$; $-1$ at $(j,j)$; $\lambda_{ji}$ at $(i,j)$; and $-\lambda_{ji}$ at $(i,i)$.
Without loss of generality, we assume $1\leq i<j\leq n$. Dividing Eq. (A.2) by $\Delta P_{ji}$ and letting $\Delta P_{ji}\to 0$, we get the following equation for the derivative of $\pi$ with respect to $P_{ji}$:
@@
where the right-hand side has all zeros except the $i$th and the $j$th elements.
@@
where $\eta_{1},\cdots,\eta_{6}$ are column vectors, $\mu_{1},\cdots,\mu_{6}$ are row vectors, and $H_{1},\cdots,H_{9}$ are sub-matrices of $P^{\top}$.
To solve $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}}$ from Eq. (A.3), we write the augmented matrix for the identity $\frac{\mathrm{d}\pi}{\mathrm{d}P_{ji}}\vec{1}_{n}=0$ and the transpose of Eq. (A.3) as
@@
Since all rows of Eq (A.5), except for the first one, add up to a zero vector, we multiply the following non-singular matrix
$\left[\begin{array}[]{c|c|c|c|c|c}1&\vec{0}_{i-1}^{\top}&0&\vec{0}_{j-i-1}^{\top}&0&\vec{0}_{n-j}^{\top}\\ \vec{0}_{i-1}&I_{i-1}&\vec{0}_{i-1}&\vec{0}_{i-1}\vec{0}_{j-i-1}^{\top}&\vec{0}_{i-1}&\vec{0}_{i-1}\vec{0}_{n-j}^{\top}\\ 0&\vec{1}_{i-1}^{\top}&\boxed{1}&\vec{1}_{j-i-1}^{\top}&\boxed{1}&\vec{1}_{n-j}^{\top}\\ \vec{0}_{j-i-1}&\vec{0}_{j-i-1}\vec{0}_{i-1}^{\top}&\vec{0}_{j-i-1}&I_{j-i-1}&\vec{0}_{j-i-1}&\vec{0}_{j-i-1}\vec{0}_{n-j}^{\top}\\ 0&\vec{0}_{i-1}^{\top}&\boxed{0}&\vec{0}_{j-i-1}^{\top}&\boxed{1}&\vec{0}_{n-j}^{\top}\\ \vec{0}_{n-j}&\vec{0}_{n-j}\vec{0}_{i-1}^{\top}&\vec{0}_{n-j}&\vec{0}_{n-j}\vec{0}_{j-i-1}^{\top}&\vec{0}_{n-j}&I_{n-j}\end{array}\right]$ (A.6)
to the left side of Eq. (A.5). Because Eq. (A.6) is the $(n+1)\times(n+1)$ identity matrix except for $n$ ones in the $(i+1)$st row, the multiplication adds the rows of Eq. (A.5) (except for the first row) to its $(i+1)$st row. Consequently, the $(i+1)$st row of Eq. (A.5) becomes a zero vector and all others remain unchanged.
After dropping the $(i+1)$st row in Eq. (A.5) and moving the $i$th column to the first column without changing the order of other columns, we obtain the augmented matrix for $\left(\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}},\frac{\mathrm{d}\pi_{-i}^{\top}}{\mathrm{d}P_{ji}}\right)^{\top}$ as
@@
Similarly, after we multiply the following non-singular matrix
$\left[\begin{array}[]{c|c|c|c|c|c}1&\vec{0}_{i-1}^{\top}&0&\vec{0}_{j-i-1}^{\top}&0&\vec{0}_{n-j}^{\top}\\ \vec{0}_{i-1}&I_{i-1}&\vec{0}_{i-1}&\vec{0}_{i-1}\vec{0}_{j-i-1}^{\top}&\vec{0}_{i-1}&\vec{0}_{i-1}\vec{0}_{n-j}^{\top}\\ 0&\vec{0}_{i-1}^{\top}&\boxed{1}&\vec{0}_{j-i-1}^{\top}&\boxed{0}&\vec{0}_{n-j}^{\top}\\ \vec{0}_{j-i-1}&\vec{0}_{j-i-1}\vec{0}_{i-1}^{\top}&\vec{0}_{j-i-1}&I_{j-i-1}&\vec{0}_{j-i-1}&\vec{0}_{j-i-1}\vec{0}_{n-j}^{\top}\\ 0&\vec{1}_{i-1}^{\top}&\boxed{1}&\vec{1}_{j-i-1}^{\top}&\boxed{1}&\vec{1}_{n-j}^{\top}\\ \vec{0}_{n-j}&\vec{0}_{n-j}\vec{0}_{i-1}^{\top}&\vec{0}_{n-j}&\vec{0}_{n-j}\vec{0}_{j-i-1}^{\top}&\vec{0}_{n-j}&I_{n-j}\end{array}\right]$
to the left side of Eq. (A.5), the $(j+1)$st row of Eq. (A.5) becomes a zero vector and all others remain unchanged. After dropping the $(j+1)$st row in Eq. (A.5) and moving the $j$th column to the first column without changing the order of other columns, we obtain the augmented matrix for $\left(\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ji}},\frac{\mathrm{d}\pi_{-j}^{\top}}{\mathrm{d}P_{ji}}\right)^{\top}$ as
@@
We next multiply the following non-singular matrix
@@
A3. Proof of Corollary 2
First, both $(I_{n-1}-Z_{i})^{-1}=I_{n-1}+Z_{i}+Z_{i}^{2}+Z_{i}^{3}+\cdots$ and $(I_{n-1}-Z_{j})^{-1}=I_{n-1}+Z_{j}+Z_{j}^{2}+Z_{j}^{3}+\cdots$ are non-negative matrices. Secondly, $\vec{1}_{n-1}$, $\gamma_{ji}$, $\gamma_{ij}$, $\alpha_{i}$, and $\alpha_{j}$ are all non-negative. Therefore, by Eq. (9) and (10) in Theorem 2, $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}}\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ji}}\leq 0$. Also,
@@
A4. Proof of Theorem 4
@@
From the first row, country $i$ could not increase $\pi_{i}$ and achieve a trade surplus with country $j$ at the same time. When $\pi_{j}/\pi_{i}<\lambda_{ji}<g_{j}/g_{i}$ in the same row, however, country $j$ has a trade surplus, an increasing $\pi_{j}$, and $\Delta P_{ji}>0$. In these ten scenarios, on the one hand, none strictly dominates the others, regarding competitiveness, net trade balance, comparative advantages, and their magnitudes of changes. On the other hand, countries $i$ and $j$ may not completely contradict in all these interests.
A5. Proof of Theorem 5
Let $x=\lambda_{ji}$. Then $\lambda_{ij}=1/x$. By Eq. (16), the Nash bargaining solution is
@@
A6. Proof of Corollary 3
When we extend the effective bargaining set of $\lambda_{ji}$ from $[p,q]$ to $[cp,q/c]$, the corresponding range for $\lambda_{ij}$ then changes from $[1/q,1/p]$ to $[c/q,1/(cp)]$. Letting $x=\lambda_{ji}$, the bargaining solution to Eq. (16) becomes
@@
After we replace $p$ and $q$ with $cp$ and $q/c$, respectively, Eq. (17) or (18) becomes
@@
A7. Proof of Theorem 6
With the change $\Delta P_{ii}$ on $P_{ii}$, we have $\Delta P=\frac{\Delta P_{ii}}{1-P_{ii}}M$ and so, $\frac{\mathrm{d}P}{\mathrm{d}P_{ii}}=\frac{M}{1-P_{ii}}$ as $\Delta P_{ii}\to 0$. We write $\frac{M}{1-P_{ii}}$ as $\frac{M}{1-P_{ii}}=[\xi_{1},\xi_{2},\cdots,\xi_{n}]$ where $\xi_{j}$ is the $j$th column of $\frac{M}{1-P_{ii}}$. Then $\xi_{1}+\xi_{2}+\cdots+\xi_{n}=\vec{0}_{n}$. We divide Eq. (A.2) by $\Delta P_{ii}$ and let $\Delta P_{ii}\to 0$ to get
$\frac{\mathrm{d}\pi}{\mathrm{d}P_{ii}}[I_{n}-P]=\pi\frac{\mathrm{d}P}{\mathrm{d}P_{ii}}=\pi\frac{M}{1-P_{ii}}=\left(\pi\xi_{1},\cdots,\pi\xi_{n}\right).$ (A.12)
With the partition Eq. (A.4) of $P^{\top}$, the augmented matrix for the identity $\frac{\mathrm{d}\pi}{\mathrm{d}P_{ii}}\vec{1}_{n}=0$ and the transpose of Eq. (A.12) is
@@
As the sum of the last column of (A.13) $\sum\limits_{j=1}^{n}\pi\xi_{j}=\pi\sum\limits_{j=1}^{n}\xi_{j}=0$, we multiply the nonsingular matrix of Eq. (A.6) to the left side of Eq. (A.13) to make the $(i+1)$st row of Eq. (A.13) to zeros. After dropping the $(i+1)$st row in Eq. (A.13) and moving the $i$th column to the first column without changing the order of other columns, we get the augmented matrix for $\left(\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}},\frac{\mathrm{d}\pi_{-i}^{\top}}{\mathrm{d}P_{ii}}\right)^{\top}$ as
$\left[\begin{array}[]{cc | c}1&\vec{1}_{n-1}^{\top}&0\\ -\alpha_{i}&I_{n-1}-Z_{i}&\frac{1}{1-P_{ii}}(\pi M_{i})^{\top}\end{array}\right].$ (A.14)
We next multiply the matrix of Eq. (A.8) to the left side of Eq. (A.14) to get
$\left[\begin{array}[]{cc | c}1+\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{i})^{-1}\alpha_{i}&\vec{0}_{n-1}^{\top}&-\frac{1}{1-P_{ii}}\vec{1}^{\top}_{n-1}(I_{n-1}-Z_{i})^{-1}(\pi M_{i})^{\top}\\ -\alpha_{i}&I_{n-1}-Z_{i}&\frac{1}{1-P_{ii}}(\pi M_{i})^{\top}\end{array}\right].$
Therefore, by the first row of Eq. (A.15),
$\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}=-\frac{1}{1-P_{ii}}\frac{\vec{1}^{\top}_{n-1}(I_{n-1}-Z_{i})^{-1}(\pi M_{i})^{\top}}{1+\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{i})^{-1}\alpha_{i}}.$
Also, by the second row of Eq. (A.15),
$-\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}\alpha_{i}+\left(I_{n-1}-Z_{i}\right)\frac{\mathrm{d}\pi_{{}_{-i}}}{\mathrm{d}P_{ii}}=\frac{1}{1-P_{ii}}(\pi M_{i})^{\top}$
and thus
$\begin{array}[]{rcl}\frac{\mathrm{d}\pi_{{}_{-i}}}{\mathrm{d}P_{ii}}&=&\left(I_{n-1}-Z_{i}\right)^{-1}\left[\frac{1}{1-P_{ii}}(\pi M_{i})^{\top}+\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}\alpha_{i}\right]\\ &=&\frac{\left(I_{n-1}-Z_{i}\right)^{-1}}{1-P_{ii}}\left[(\pi M_{i})^{\top}-\frac{\vec{1}^{\top}_{n-1}(I_{n-1}-Z_{i})^{-1}\left(\pi M_{i}\right)^{\top}}{1+\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{i})^{-1}\alpha_{i}}\alpha_{i}\right].\end{array}$
A8. Proof of Corollary 4
Part 1: If $\pi M_{i}=\vec{0}_{n-1}^{\top}$, then $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}=0$ and $\frac{\mathrm{d}\pi_{-i}}{\mathrm{d}P_{ii}}=\vec{0}_{n-1}$ according to Eq. (22) and Eq. (23). Thus, $\frac{\mathrm{d}\pi}{\mathrm{d}P_{ii}}=\vec{0}_{n}^{\top}$.
Conversely, if $\frac{\mathrm{d}\pi}{\mathrm{d}P_{ii}}=\vec{0}_{n}^{\top}$, then $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}=0$ and $\frac{\mathrm{d}\pi_{-i}}{\mathrm{d}P_{ii}}=\vec{0}_{n-1}$. Applying $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}=0$ to Eq. (22), we have
@@
We plug Eq. (A.16) into Eq. (23) and use $\frac{\mathrm{d}\pi_{-i}}{\mathrm{d}P_{ii}}=\vec{0}_{n-1}$ to get
$\frac{\left(I_{n-1}-Z_{i}\right)^{-1}}{1-P_{ii}}\left(\pi M_{i}\right)^{\top}=\vec{0}_{n-1}.$
Therefore, $\left(\pi M_{i}\right)^{\top}=(1-P_{ii})\left(I_{n-1}-Z_{i}\right)\vec{0}_{n-1}=\vec{0}_{n-1}$ and $\pi M_{i}=\vec{0}_{n-1}^{\top}.$
Part 2: Using the definition of $M_{i}$, $\pi M_{i}=\vec{0}_{n-1}^{\top}$ implies that
$\sum\limits_{j\not=i}\frac{\lambda_{ij}P_{ij}P_{jt}}{1-P_{ji}}\pi_{j}-\pi_{i}P_{it}=0$
for all $t\not=i$. We sum the above equation over all $t\not=i$ to get
@@
which leads to
$\sum\limits_{j\not=i}\lambda_{ij}P_{ij}\pi_{j}=\pi_{i}-\pi_{i}P_{ii}$ (A.17)
and
$\sum\limits_{j=1}^{n}\pi_{j}\left[\Lambda\bigotimes P\right]_{ji}^{\top}=\pi_{i}.$
When $i$ runs all numbers in $\mathcal{N}$, thus, $\pi\left[\Lambda\bigotimes P\right]^{\top}=\pi$.
Finally, if $\lambda_{ij}=\pi_{i}/\pi_{j}$, then Eq. (A.17) is clearly valid and so is Eq. (24). If $\lambda_{ij}=P_{ji}/P_{ij}$, then Eq. (4) implies Eq. (A.17) and, therefore, Eq. (24) also holds.

### v5

@@ 7 Discussions
We let $P_{ji}$ have a small change $\Delta P_{ji}$ and attempt to calculate its effect on $\pi$, including $\pi_{i}$, $\pi_{-i}$, and $\pi_{j}$. Accordingly, $P$ has three other simultaneous changes: $-\Delta P_{ji}$ on $P_{jj}$ to offset the change on $P_{ji}$ in the $j$th row; $\lambda_{ji}\Delta P_{ji}$ on $P_{ij}$ for country $j$’s retaliation upon $i$’s change at $P_{ji}$; and $-\lambda_{ji}\Delta P_{ji}$ on $P_{ii}$ to maintain the unit sum of the $i$th row. Therefore, $\frac{\mathrm{d}P}{\mathrm{d}P_{ji}}=\lim\limits_{\Delta P_{ji}\to 0}\frac{\Delta P}{\Delta P_{ji}}$ is a zero $n\times n$ matrix except for: $1$ at $(j,i)$; $-1$ at $(j,j)$; $\lambda_{ji}$ at $(i,j)$; and $-\lambda_{ji}$ at $(i,i)$. Without loss of generality, we assume $1\leq i<j\leq n$. Dividing Eq. (A.2) by $\Delta P_{ji}$ and letting $\Delta P_{ji}\to 0$, we get the following equation for the derivative of $\pi$ with respect to $P_{ji}$:
@@
where the right-hand vector has all zeros except for the $i$th and the $j$th elements.
@@
where $\eta_{1},\cdots,\eta_{6}$ are column vectors, $\mu_{1},\cdots,\mu_{6}$ are row vectors, and $H_{1},\cdots,H_{9}$ are sub-matrices of $P^{\top}$. We write the augmented matrix for the identity $\frac{\mathrm{d}\pi}{\mathrm{d}P_{ji}}\vec{1}_{n}=0$ and the transpose of Eq. (A.3) as
@@
Since all rows of Eq (A.5), except for the first one, sum up to a zero vector, we add them to the $(i+1)$st row, making the $(i+1)$st row of Eq. (A.5) a zero vector. After dropping the $(i+1)$st row in Eq. (A.5) and moving the $i$th column to the first column without changing the order of other columns, we obtain the augmented matrix for $\left(\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}},\frac{\mathrm{d}\pi_{-i}^{\top}}{\mathrm{d}P_{ji}}\right)^{\top}$:
@@
Similarly, after we add all rows (except for the first one) in Eq. (A.5) to the $(j+1)$st row, the $(j+1)$st row becomes a zero vector. After dropping the $(j+1)$st row in Eq. (A.5) and moving the $j$th column to the first column without changing the order of other columns, we obtain the augmented matrix for $\left(\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ji}},\frac{\mathrm{d}\pi_{-j}^{\top}}{\mathrm{d}P_{ji}}\right)^{\top}$:
@@
We multiply the following non-singular matrix:
@@
A4. Proof of Corollary 3
First, both $(I_{n-1}-Z_{i})^{-1}=I_{n-1}+Z_{i}+Z_{i}^{2}+Z_{i}^{3}+\cdots$ and $(I_{n-1}-Z_{j})^{-1}=I_{n-1}+Z_{j}+Z_{j}^{2}+Z_{j}^{3}+\cdots$ are non-negative matrices. Secondly, $\vec{1}_{n-1}$, $\gamma_{ji}$, $\gamma_{ij}$, $\alpha_{i}$, and $\alpha_{j}$ are all non-negative. Therefore, by Eqs. (9) and (10) in Theorem 2, $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}}\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ji}}\leq 0$. Also,
@@
A5. Proof of Theorem 3
In added value, the changed exports from country $j$ to $i$ are $g_{j}\Delta P_{ji}$, while those from country $i$ to $j$ are $g_{i}\Delta P_{ij}=g_{i}\lambda_{ji}\Delta P_{ji}$. If $\lambda_{ji}=g_{j}/g_{i}$, then $g_{j}\Delta P_{ji}=g_{i}\Delta P_{ij}$, resulting in zero net trade surplus and deficit between these countries. This condition is for instant balance change.
Over a specific time period, there is no trade surplus or deficit if and only if $g_{j}/g_{i}=P_{ij}/P_{ji}$, because $g_{i}P_{ij}$ represents exports from $i$ to $j$ and $g_{j}P_{ji}$ represents exports from $j$ to $i$.
A6. Proof of Theorem 4
@@
From the first row, country $i$ could not increase $\pi_{i}$ and achieve a trade surplus with country $j$ at the same time. When $\pi_{j}/\pi_{i}<\lambda_{ji}<g_{j}/g_{i}$ in the same row, however, country $j$ has a trade surplus, an increasing $\pi_{j}$, and $\Delta P_{ji}>0$. In these ten scenarios, none strictly dominates the others, regarding competitiveness, net trade balance, comparative advantages, and their magnitudes of changes. On the other hand, countries $i$ and $j$ may not completely contradict in all these interests.
A7. Proof of Theorem 5
In the first scenario, country $j$ prefers $q$ to $p$, whereas country $i$ prefers $1/p$ to $1/q$. Therefore, $u_{j}(x)=x$ is a utility function for $j$ because $u_{j}(q)>u_{j}(p)$. Similarly, $u_{i}(x)=x$ is a utility function for $i$. These are von Neumann-Morgenstern utilities, uniquely determined up to a positive affine transformation. Assuming that all countries have equal prior bargaining power, the Nash bargaining solution solves the following maximization problem:
$\max\limits_{\lambda_{ji}\lambda_{ij}=1;\ \lambda_{ji}>0}\left(\lambda_{ji}-p\right)\left(\lambda_{ij}-\frac{1}{q}\right).$ (A.11)
In Eq. (A.11), $\lambda_{ji}-p$ and $\lambda_{ij}-\frac{1}{q}$ are the excessive payoffs or welfare for countries $j$ and $i$ in terms of their respective utility functions. The bargaining solution maximizes the product of the excessive utilities.
In the second scenario of opposite preferences, country $j$ prefers $p$ to $q$ while country $i$ prefers $1/q$ to $1/p$. Their utility functions could be $u_{j}(x)=u_{i}(x)=-x$, and their status quo points are $u_{j}(q)=-q$ and $u_{i}(1/p)=-1/p$, respectively. The maximization problem becomes:
$\max\limits_{\lambda_{ji}\lambda_{ij}=1;\ \lambda_{ji}>0}\left(-\lambda_{ji}+q\right)\left(-\lambda_{ij}+\frac{1}{p}\right).$ (A.12)
In this utility function, $\lambda_{ji}$ seems to have no meaning of reprisal. However, we can rephrase Eq. (A.12) by:
$\max\limits_{\lambda_{ji}\lambda_{ij}=1;\ \lambda_{ji}>0}\left(\lambda_{ji}-q\right)\left(\lambda_{ij}-\frac{1}{p}\right)$ (A.13)
in which $\lambda_{ji}-q$ and $\lambda_{ij}-1/p$ are the loss functions for $j$ and $i$, respectively. To coerce an agreement formation, each country plays a threat strategy to maximize its counterpart’s loss. The final outcome is Eq. (A.13). Moving from $p$ to $q$, a larger $\lambda_{ji}$ means less loss to $j$; thus, $\lambda_{ji}$ still preserves the meaning of retaliation for $j$ in Eq. (A.13) and, so, in Eq. (A.12).
Let $x=\lambda_{ji}$. Then $\lambda_{ij}=1/x$. By Eq. (A.11), the Nash bargaining solution is:
@@
A8. Proof of Corollary 4
When we extend the effective bargaining set of $\lambda_{ji}$ from $[p,q]$ to $[cp,q/c]$, the corresponding range for $\lambda_{ij}$ then changes from $[1/q,1/p]$ to $[c/q,1/(cp)]$. Letting $x=\lambda_{ji}$, the bargaining solution to Eq. (A.11) becomes:
@@
After we replace $p$ and $q$ with $cp$ and $q/c$, respectively, Eq. (A.12) or (A.13) becomes:
@@
A9. Proof of Corollary 5
This is because:
$\underset{\lambda_{ji}>0}{\mathrm{argmax}}\left\{\left(\lambda_{ji}-\frac{\pi_{j}}{\pi_{i}}\right)\left(\lambda_{ij}-\frac{\pi_{i}}{\pi_{j}}\right)\bigg{|}\lambda_{ij}\lambda_{ji}=1\right\}=\frac{\pi_{j}}{\pi_{i}}$ (A.14)
and
$\underset{\lambda_{ji}>0}{\mathrm{argmax}}\left\{\left(\lambda_{ji}-\frac{g_{j}}{g_{i}}\right)\left(\lambda_{ij}-\frac{g_{i}}{g_{j}}\right)\bigg{|}\lambda_{ij}\lambda_{ji}=1\right\}=\frac{g_{j}}{g_{i}}.$ (A.15)
In Eqs. (A.14) and (A.15), both countries battle for better competitiveness and net trade balance, respectively. At the solution points, the objective functions are zero and no country gets extra welfare,
A10. Proof of Corollary 6
Country $j$’s direct competition with $i$ for a higher $\pi_{j}$ results in $\lambda_{ji}=\pi_{j}/\pi_{i}$. If it seeks a better trade balance instead, then the Nash solution expects $\lambda_{ji}=\sqrt{\pi_{j}g_{j}}/\sqrt{\pi_{i}g_{i}}$, which is likely larger than $\pi_{j}/\pi_{i}$ according to the Matthew effect. Consequently, by Eq. (10), $\pi_{j}$ rises as $j$’s exports to $i$ grow, regardless of trade surplus or deficit.
Similarly, if country $j$ competes with $i$ for a better trade balance, then the solution $\lambda_{ji}=g_{j}/g_{i}>\pi_{j}/\pi_{i}$ and a decreasing $P_{ji}$ leads to a declining $\pi_{j}$ in Eq. (10). Thus, $j$ should also not compete with $i$.
A11. Proof of Theorem 6
We write $M=[\xi_{1},\xi_{2},\cdots,\xi_{n}]$ where $\xi_{j}$ is the $j$th column of $M$. Since $\sigma_{ji}$ has a zero sum for all $j$, $\xi_{1}+\xi_{2}+\cdots+\xi_{n}=\vec{0}_{n}$. We divide Eq. (A.2) by $\Delta P_{ii}$ and let $\Delta P_{ii}\to 0$ to get:
$\frac{\mathrm{d}\pi}{\mathrm{d}P_{ii}}[I_{n}-P]=\pi\frac{\mathrm{d}P}{\mathrm{d}P_{ii}}=\pi M=\left(\pi\xi_{1},\cdots,\pi\xi_{n}\right).$ (A.16)
With the partition Eq. (A.4) of $P^{\top}$, the augmented matrix for the identity $\frac{\mathrm{d}\pi}{\mathrm{d}P_{ii}}\vec{1}_{n}=0$ and the transpose of Eq. (A.16) is:
@@
As the last column of Eq. (A.17) sums to zero, i.e., $\sum\limits_{j=1}^{n}\pi\xi_{j}=\pi\sum\limits_{j=1}^{n}\xi_{j}=0$, we add all rows (except for the first) to the $(i+1)$st row to make the $(i+1)$st row a zero vector. After dropping the $(i+1)$st row and moving the $i$th column to the first column without changing the order of other columns, we get the augmented matrix for $\left(\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}},\frac{\mathrm{d}\pi_{-i}^{\top}}{\mathrm{d}P_{ii}}\right)^{\top}$:
$\left[\begin{array}[]{cc | c}1&\vec{1}_{n-1}^{\top}&0\\ -\alpha_{i}&I_{n-1}-Z_{i}&(\pi M_{i})^{\top}\end{array}\right].$ (A.18)
We multiply the matrix of Eq. (A.7) to the left side of Eq. (A.18) to get:
$\left[\begin{array}[]{cc | c}1+\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{i})^{-1}\alpha_{i}&\vec{0}_{n-1}^{\top}&-\vec{1}^{\top}_{n-1}(I_{n-1}-Z_{i})^{-1}(\pi M_{i})^{\top}\\ -\alpha_{i}&I_{n-1}-Z_{i}&(\pi M_{i})^{\top}\end{array}\right].$
Therefore, by the first row of Eq. (A.19):
$\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}=-\frac{\vec{1}^{\top}_{n-1}(I_{n-1}-Z_{i})^{-1}(\pi M_{i})^{\top}}{1+\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{i})^{-1}\alpha_{i}}.$
Also, by the second row of Eq. (A.19):
$-\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}\alpha_{i}+\left(I_{n-1}-Z_{i}\right)\frac{\mathrm{d}\pi_{{}_{-i}}}{\mathrm{d}P_{ii}}=(\pi M_{i})^{\top}$
and thus:
$\begin{array}[]{rcl}\frac{\mathrm{d}\pi_{{}_{-i}}}{\mathrm{d}P_{ii}}&=&\left(I_{n-1}-Z_{i}\right)^{-1}\left[(\pi M_{i})^{\top}+\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}\alpha_{i}\right]\\ &=&\left(I_{n-1}-Z_{i}\right)^{-1}\left[(\pi M_{i})^{\top}-\frac{\vec{1}^{\top}_{n-1}(I_{n-1}-Z_{i})^{-1}\left(\pi M_{i}\right)^{\top}}{1+\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{i})^{-1}\alpha_{i}}\alpha_{i}\right].\end{array}$
A12. Proof of Corollary 7
If $\pi M_{i}=\vec{0}_{n-1}^{\top}$, then $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}=0$ and $\frac{\mathrm{d}\pi_{-i}}{\mathrm{d}P_{ii}}=\vec{0}_{n-1}$ according to Eqs. (17) and (18). Thus, $\frac{\mathrm{d}\pi}{\mathrm{d}P_{ii}}=\vec{0}_{n}^{\top}$.
Conversely, if $\frac{\mathrm{d}\pi}{\mathrm{d}P_{ii}}=\vec{0}_{n}^{\top}$, then $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}=0$ and $\frac{\mathrm{d}\pi_{-i}}{\mathrm{d}P_{ii}}=\vec{0}_{n-1}$. Applying $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}=0$ to Eq. (17), we have:
@@
We plug Eq. (A.20) into Eq. (18) and use $\frac{\mathrm{d}\pi_{-i}}{\mathrm{d}P_{ii}}=\vec{0}_{n-1}$ to get:
$\left(I_{n-1}-Z_{i}\right)^{-1}\left(\pi M_{i}\right)^{\top}=\vec{0}_{n-1}.$
Therefore, $\left(\pi M_{i}\right)^{\top}=\left(I_{n-1}-Z_{i}\right)\vec{0}_{n-1}=\vec{0}_{n-1}$ and $\pi M_{i}=\vec{0}_{n-1}^{\top}.$
A13. Proof of Corollary 8
Since
@@
$\pi M_{i}=\vec{0}_{n-1}^{\top}$ implies that $-(1-P_{ii})\pi M$, i.e.,
$\left(\pi_{1}\lambda_{i1}P_{i1},\cdots,\pi_{i-1}\lambda_{i,i-1}P_{i,i-1},-\pi_{i}(1-P_{ii}),\pi_{i+1}\lambda_{i,i+1}P_{i,i+1},\cdots,\pi_{n}\lambda_{in}P_{in}\right)\left[\begin{array}[]{c}\sigma_{1i}\\ \vdots\\ \sigma_{ni}\end{array}\right],$
is a zero vector except for its $i$th element. Thus, for any $t\neq i$, the zero value in the $t$th element implies that:
$\sum\limits_{j\neq i}\frac{\lambda_{ij}P_{ij}P_{jt}}{1-P_{ji}}\pi_{j}=\pi_{i}P_{it}.$
We sum the above equations over all $t\neq i$ to get:
$\sum_{t\neq i}\sum\limits_{j\neq i}\frac{\lambda_{ij}P_{ij}P_{jt}}{1-P_{ji}}\pi_{j}=\pi_{i}\sum_{t\neq i}P_{it},$
which leads to:
$\sum\limits_{j\neq i}\lambda_{ij}P_{ij}\pi_{j}=\pi_{i}-\pi_{i}P_{ii}$ (A.21)
and:
$\sum\limits_{j=1}^{n}\pi_{j}\left[\Lambda\odot P\right]_{ji}^{\top}=\pi_{i}.$
When $i$ runs all numbers in $\mathcal{N}$, thus, $\pi\left[\Lambda\odot P\right]^{\top}=\pi$.
Finally, if $\lambda_{ij}=\pi_{i}/\pi_{j}$, then Eq. (A.21) is clearly valid and so is Eq. (19). If $\lambda_{ij}=P_{ji}/P_{ij}$, then $\left[\Lambda\odot P\right]^{\top}=P$ and, therefore, Eq. (19) also holds.